OpenStudy (anonymous):

The sequence {Aₓ} is defined by A₁ = 2 and Aₓ₊₁ = Aₓ + 2x. What is the value of A₁₀₀. a.9902 b.9900 c. 10100 d.. 9904

OpenStudy (anonymous):

u have a1=2 a2=2+2 a3=4+4 a4=8+6 a5=14+8 a100= a99+2X99 look for the 1st term on RHS

OpenStudy (anonymous):

do u find any pattern?

OpenStudy (anonymous):

2,4,8,14,22,32....

OpenStudy (anonymous):

do u find any pattern in this sequence?

OpenStudy (anonymous):

no thats why i am not getting it

OpenStudy (anonymous):

see the difference of consecutive terms...

OpenStudy (anonymous):

yes its 2 4 6 8 10

OpenStudy (anonymous):

so these r in AP...right?

OpenStudy (anonymous):

yes so

OpenStudy (anonymous):

s= 2+4+8+14+22+32+............+a99 s= 2+4+ 8 +14+22+32+......+a98+a99

OpenStudy (anonymous):

subtract these two

OpenStudy (anonymous):

0= 2+2+4+6+..........+(a99-a98) -a99

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so a99= 2+2+4+6+..........+(a99-a98)

OpenStudy (anonymous):

now u can find a99?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

since RHS is an AP so u have to find the sum of an AP for getting the value of a99

OpenStudy (anonymous):

2+2+4+6+..........+(a99-a98) a99=2+(2+4+6+ ....98 terms)

OpenStudy (anonymous):

simply put value of x =99 that is a100 = a99+198. now u can find an a99 by using an formula of AP on series of a1, a2,a3,a4,a5. and get a99 and put tha value of a99 in a100 and get your result.

OpenStudy (anonymous):

a99= 2+ 98/2{ 2*2 +(98-1)2}

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):

a99=9800

OpenStudy (anonymous):

check it

OpenStudy (anonymous):

no ..there was a mistake

OpenStudy (raden):

it should be : a100 = 2 + {s99} = 2 + 9900

OpenStudy (anonymous):

yes it will be 9902

OpenStudy (raden):

with s99 = 99/2 (2+198)